OpenStudy (anonymous):

Solve: sqrt{3x-5} + sqrt{2x+3} +1 =0

OpenStudy (cwrw238):

3x + 5 + 2x + 3 + 2sqrt(3x+5)(2x+3) = -1 2sqrt(3x+5)(2x+3 = 5x - 9 square both sides 4(3x+5(2x+3) = (5x - 9)^2 this will reduce to a quadratic equation

OpenStudy (anonymous):

still confused.... sorry

OpenStudy (cwrw238):

expand both sides 24x^2 + 76x + 60 = 25x + 90x + 81 x^2 + 14x + 21 = 0

OpenStudy (anonymous):

ok, thank you for the information.. I really am only an English tutor trying to help math students

OpenStudy (anonymous):

so the answer would be x^2 + 14x + 21 = 0 ..??

OpenStudy (cwrw238):

however with these type of problems because of the squaring you cannot be sure that these solutions are valid - in fact these 2 are not - if you plug them into original equation they will not fit because you cant have a real square root of a negative number eg sqrt (-1.71*3 -5) there are no solutions to this equation - the best way would have been to draw the graph i'll do that on wolfram

OpenStudy (cwrw238):

here it is ://www.wolframalpha.com/input/?i=+sqrt%7B3x5%7D+%2B+sqrt%7B2x%2B3%7D+%2B1+%3D0

OpenStudy (anonymous):

ok, I appreciate that very much!

OpenStudy (cwrw238):

the graph looks like this 

OpenStudy (cwrw238):

if there were any roots the curve would intersect the x-axis yw

OpenStudy (anonymous):

ok, thank you.. GOT IT!!

OpenStudy (anonymous):

YOU ARE SO AWESOME!!

OpenStudy (cwrw238):

lol - ty - but it would have been wiser to draw the graph first!